OpenStudy (anonymous):

1.For the following questions: solve each proportion for the appropriate letter. a) m + 12/9m = 5/9 b) 12/x – 2 = 24/3x + 3 c)y/6 = y – 4/2 Medals will be rewarded @tHe_FiZiCx99

OpenStudy (anonymous):

any idea

OpenStudy (anonymous):

\[M + 12 \over 9m\]

OpenStudy (anonymous):

= \[ 5 \over 9\]

OpenStudy (the_fizicx99):

Hold on, typo.

OpenStudy (anonymous):

ok

OpenStudy (the_fizicx99):

m + 12/9m = 5/9 9(m + 12) = 5(9m) cross multiply m + 12 = 5m divide ( 9's cancel out) then subtract both sides m 12 = 4m subtract m = 3 divide to solve for m

OpenStudy (anonymous):

So,

OpenStudy (anonymous):

is 4 the slope

OpenStudy (the_fizicx99):

What? No. This is a different problem

OpenStudy (anonymous):

I'm so confused right now

OpenStudy (the_fizicx99):

This isn't the same problem as before. This is proportions. Not slopes. I gtg, can you do #2 and #3 now? Use mine as a guide :P

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

help me @csscgirl

OpenStudy (the_fizicx99):

Ok! Gl :p

undeadknight26 (undeadknight26):

@jdoe0001 @shamil98

OpenStudy (anonymous):

@jdoe0001 and @shamil98 can you help me